undeadknight26 (undeadknight26):

Given a polynomial f(x), if (x + 2) is a factor, what else must be true? f(0) = 2 f(0) = -2 f(2) = 0 f(-2) = 0

undeadknight26 (undeadknight26):

Im pretty sure is D...Just checking though.

OpenStudy (cwrw238):

yes its the factor theorem which states that if (x - a) is a factor of a polynomial then f(a) = 0.

OpenStudy (cwrw238):

so (x + 2) = (x - (-2)) thus f(-2) = 0

OpenStudy (solomonzelman):

you are correct. because call any (second) factor (x+k) you have: f(x)=(x-2)(x+k) and the zero product property gives you that when one of the factors is equal to zero, the f(x)=0. if x=2, then (x-2) is equal to 0.........

OpenStudy (solomonzelman):

btw, check the feedback section.